OpenStudy (anonymous):

Please Help! what are the zeros of the function? what are their multiplicities? f(x)=x^4-4x^3+3x^2

jimthompson5910 (jim_thompson5910):

hint: f(x)=x^4-4x^3+3x^2 0 = x^4-4x^3+3x^2 x^4-4x^3+3x^2 = 0 x^2(x^2-4x+3) = 0 x^2(x-3)(x-1) = 0

OpenStudy (anonymous):

I am still really confused I dont get what the question is asking me!

jimthompson5910 (jim_thompson5910):

it's asking you to solve 0 = x^4-4x^3+3x^2 for x

jimthompson5910 (jim_thompson5910):

which is equivalent to x^2(x-3)(x-1) = 0

OpenStudy (anonymous):

it wants to know if its A. The numbers -1 and -3 are zeroes of multiplicities2;the numbers 1 and 3 are zeroes of multiplicities 1. B. the number 0 is a zero of multiplicity2;the numbers 1 and 3 are zeroes of multiplicty 1. C. The numbers 0 and 1 are the zeroes of multpilcty2; the number 3 is a zero of multiplcty1. D.The number 0 is zero of multiplicty2;the numbers -1 and -3 are zeroes of multiplicty

jimthompson5910 (jim_thompson5910):

x^2(x-3)(x-1) = 0 x^2 = 0, x-3=0, or x-1=0

jimthompson5910 (jim_thompson5910):

solve each for x

OpenStudy (anonymous):

x=-x^2 x=3 x=1

OpenStudy (anonymous):

so how do i know which one it is?

jimthompson5910 (jim_thompson5910):

x^2 = 0 ---> x = 0 (multiplicity of 2)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks so much!! I really apperciate it!

jimthompson5910 (jim_thompson5910):

np